OpenStudy (anonymous):

Solve 2 log x = log 64. x = 8 x = ±8 x = 32 x = 128

OpenStudy (anonymous):

i think its b but im not sure b is plus or minus 8

OpenStudy (anonymous):

2 \[2*32=64\]

OpenStudy (mathmale):

Please, would you show your work and mention which of the rules of logarithms you're using? Note that if you arrive at a particular answer, you can and should check it yourself by substituting the x value into the given equation. Note also that -8 could not be a solution, because the domain of y = ln x is "all real numbers GREATER than zero."

OpenStudy (mathmale):

Rules of logs: 1. log of a product: ln x*y 2. log of a quotient: ln (x/y) 3. log of a power: ln (x^y)

OpenStudy (anonymous):

math male was my answer correct

OpenStudy (anonymous):

\[ \log _{a}(x^p) = p \log _{a} x \] \[ \log (x^2) = 2 \log x \]

OpenStudy (mathmale):

Purdy: Please see my comments and suggestions, above, especially that involving x=-8.

OpenStudy (mathmale):

sangya21's suggestion is a very good one. Can you apply it?

OpenStudy (anonymous):

Yes, thank you i got the answer

OpenStudy (anonymous):

Simplify 5 log2 k − 8 log2 m + 10 log2 n. 7 log2 (k − m + n) 7 log2 kn over m log2 50 kn over 8 m log2 k to the fifth power n to the tenth power over m to the eighth power

OpenStudy (anonymous):

can u help me with this question also?

OpenStudy (mathmale):

Great. If you have another math problem you want to discuss, kindly enter it as a NEW question.

OpenStudy (anonymous):

do i have to type the whole thing if it dosen't paste?

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

Simplify 5 log2 k-8 log2 m+10 log2 n